# Semantic Mapping in Indoor Embodied AI – A Comprehensive Survey and Future Directions

### v1

@@ 6 Map Evaluation
In the embodied AI literature, very little focus has been given to evaluating the built map. Instead the focus has been to evaluate the agent performance on the downstream tasks through various metrics. However, it is crucial to evaluate the semantic maps in terms of accuracy, completeness, consistency and robustness in addition to its utility on the downstream tasks. We next discuss these evaluation metrics in detail.
Utility. Most works build semantic maps as an intermediate step while performing downstream tasks, such as navigation, exploration, manipulation and so on. In such cases, the map is exploited by the task planner to plan a path and generate low-level actions in order to complete the task. Moreover, most of these works focus on a single task and hence it suffices for them to evaluate the task performance directly without caring about how well the map representation is. Gupta et al. (2017); Kim et al. (2023) build map for navigation tasks, while Chaplot et al. (2019) builds map to efficiently explore the environment. Although the open-vocabulary map in Gu et al. (2023) is pre-built once and used in multiple downstream tasks (navigation, manipulation, object segmentation etc.), they still evaluate the task performance and not the map itself. The navigation tasks are evaluated on Success and SPL (Anderson et al., 2018a), while the exploration tasks are evaluated on the coverage area (Cov) and the percentage of area explored (Cov) (Chaplot et al., 2019).
@@
Accuracy. Map accuracy refers to how accurately the map captures semantic information when compared against the ground truth. However obtaining the ground truth map can be challenging in most cases. Cartillier et al. (2021) and Georgakis et al. (2022b) evaluate the accuracy of the built 2D semantic map using semantic segmentation metrics such as pixel-wise labeling accuracy, pixel-based F1 score, Intersection-over-Union (IoU) score and a contour-based average boundary F1 score. While evaluating accuracy for semantic maps that store a fixed set of object categories is straightforward, it could be tricky to evaluate semantic maps that store implicit features. For example, OpenScene (Peng et al., 2023) performs semantic segmentation on their implicit map for a fixed set of categories and then evaluate it using accuracy and IoU. ConceptGraphs (Gu et al., 2023), on the other hand, evaluate their open-vocabulary topological map by employing human evaluators on Amazon Mechanical Turk (AMT) to report the scene graph accuracy.
@@
Completeness. Map completeness measures how completely a generated map represents an environment, encompassing both geometric and semantic coverage. Geometric coverage refers to the fraction of environment that has been mapped whereas semantic coverage refers to the completeness of the semantic information captured in the map. Completeness is particularly critical for tasks like search and rescue, where a thorough understanding of the surroundings is essential. Moreover, complete geometric and semantic mapping support better decision-making by reducing dependence on incomplete or inaccurate data. The extent of map completeness depends on how thoroughly the robot explores the environment during the downstream task and is closely tied to the ‘stopping criteria’ – a method that determines when to end the exploration process. In embodied AI, exploration typically ends when the task is deemed complete or when a predefined time budget is reached. However, reaching this time budget does not always guarantee that the map is fully complete, making the development of reliable stopping criteria an ongoing challenge in robotics as well as embodied AI research (Placed & Castellanos, 2022; Luperto et al., 2024). Geometric coverage metric has been reported in prior works that performs the task of exploration (Chaplot et al., 2019) in terms of the fraction of the environment explored by the robot. However, a lot of works that tackle other embodied AI tasks fail to report this metric even if exploration is a crucial part of their task (Gervet et al., 2023; Raychaudhuri et al., 2023; Yokoyama et al., 2023). Semantic Coverage on the other hand relies on the presence of a detailed ground-truth map with semantic information and could be hard to obtain, as discussed in the previous paragraph ‘Accuracy’. Although some works (Cartillier et al., 2021; Georgakis et al., 2022b) report semantic accuracy, none report semantic coverage of the built map.
@@
Consistency. Geometric consistency of a map refers to how accurately the spatial structure (distances, angles, and relative positions of structures/objects) of the map represents the physical layout of the environment. Accurate geometry ensures safe and efficient path planning and obstacle avoidance. This is crucial in classical robotics for identifying loop closure, a popular technique to recognize previously visited locations in SLAM, which refines the map and reduces drift. However, in embodied AI systems, there is no drift due to absence of sensory and actuator noise and hence the generated maps are mostly consistent with the environment. Hence prior embodied AI works do not report geometric consistency metrics for the built map. That said, ensuring geometric consistency will be crucial in future where dynamic moving objects may disrupt the map’s structural fidelity. It can be measured by reporting the Root Mean Square Error (RMSE) from the ground-truth geometry. On the other hand, the semantic consistency of a map refers to the alignment between the semantic information of structures/objects and their physical locations in the environment. Semantic consistency is particularly crucial to remain consistent over time despite changes in perspective, lighting, or environment dynamics when the robot moves around the environment. This could be measured using a temporal accuracy metric that measures how accuracy of the semantic information changes over time. However, none of the prior works measure semantic consistency and may be considered in future research.
Robustness. Evaluating robustness in semantic maps is mostly crucial for assessing their reliability in unpredictable or dynamic environments. A robust map exhibits low uncertainty and high confidence in its semantic information, allowing the robot to adapt to errors, sensor noise, and environmental changes. Since recent approaches use pretrained models, measuring the confidence of the model’s predictions could be beneficial to assess map robustness, with higher confidence indicating a more robust map. Alternatively, model uncertainty can be measured by assessing the variance in model predictions, reflecting epistemic uncertainty in the model, where low uncertainty will indicate a robust map. While some studies (Georgakis et al., 2022a; Raychaudhuri et al., 2024) incorporate uncertainty into task planning, it has not been used as a formal metric in previous research, presenting an area for future exploration.
In summary, there is considerable scope for future research to develop better metrics that evaluate the built semantic map and not just the downstream task performance. This can further be improved by developing a standardized evaluation framework for semantic maps.

### v2

@@ 6 Evaluation
In the embodied AI and robotics literature, very little focus has been given to evaluating the built map (intrinsic evaluation). Instead the focus has been to evaluate the agent performance on the downstream tasks (extrinsic evaluation) through various metrics. This is because these works build semantic maps as an intermediate step while performing downstream tasks, such as localization, navigation, exploration, manipulation and so on. Moreover, most of these works focus on a single task and hence it suffices for them to evaluate the task performance directly without caring about how well the map representation is. However, we think that it is crucial to do intrinsic evaluation based on the accuracy, completeness, consistency and robustness of the built map, in addition to evaluating the task performance. We next discuss the evaluation metrics (both extrinsic and intrinsic) in detail.
6.1 Extrinsic evaluation
This section discusses extrinsic (or task-level) evaluation metrics commonly used to assess agent performance across embodied tasks. In navigation, success is not simply reaching the goal but signaling it intentionally, often via a dedicated action like ‘Stop’ or ‘Found’ (Anderson et al., 2018a). Accordingly, Success Rate (SR) is used to evaluate whether the agent correctly signals goal completion, while Success weighted by inverse Path Length (SPL) accounts for trajectory efficiency. These metrics are standard in navigation benchmarks (Beeching et al., 2020; Batra et al., 2020b; Chaplot et al., 2020a; b; Wijmans et al., 2019b; Li et al., 2021c; Krantz et al., 2022; Georgakis et al., 2022a; Taioli et al., 2024; Yokoyama et al., 2024). Additional metrics include Oracle SR (OSR), which accounts for when the agent reaches the goal but fails to signal, and Navigation Error (NE), which measures the final distance to the goal. In multi-step tasks or multi-goal navigation, metrics like Progress (Wani et al., 2020; Khanna et al., 2024) or Goal-Condition Success (Shridhar et al., 2020; Padmakumar et al., 2022) capture the proportion of successfully completed subgoals. For instruction following, normalized Dynamic Time Warping (nDTW) is used to evaluate how well the agent’s trajectory aligns with the reference path (Ilharco et al., 2019; Krantz et al., 2020; Raychaudhuri et al., 2021; 2025). Tasks involving classification or detection often use precision (how often predictions are correct), recall (how well all relevant targets are found), F1 score (combines and balances precision and recall) and receiver-operating characteristic curve or ROC curve (measures the tradeoff between sensitivity or true positive rate and specificity or false positive rate) (Chen & Mooney, 2011a; Magnusson et al., 2009; Matuszek et al., 2012; Tellex et al., 2011b; Wellhausen et al., 2020) to evaluate prediction quality. In exploration tasks, coverage is used to measure the proportion of the environment observed (Chaplot et al., 2019). In manipulation tasks such as object rearrangement or tabletop operations, task success rate, measuring the percentage of trials where the agent achieves the desired manipulation (e.g., grasping, placing), is the dominant metric (Kalashnikov et al., 2018; Batra et al., 2020a; Lin et al., 2020; Zeng et al., 2020; Huang et al., 2023b; Gu et al., 2023a). Some works also evaluate policy outputs using precision (Zeng et al., 2017), or employ pose error metrics to assess deviations between the final and target object positions (Liu et al., 2012).
Summary. Extrinsic evaluation metrics have been extensively studied across the literature (Deitke et al., 2022) and are continually evolving in response to the growing complexity and diversity of embodied tasks. As new challenges emerge in navigation, manipulation, and instruction following, the need for nuanced and task-specific performance metrics has driven the development of more robust and comprehensive evaluation frameworks. Although extrinsic metrics have dominated evaluation in the field, we believe it’s time for the community to place greater emphasis on intrinsic metrics (we discuss this in next section). This shift is especially important given the recent progress in building open-vocabulary, flexible, and general-purpose maps, which require more nuanced assessment of map quality beyond task success alone.
6.2 Intrinsic evaluation
In this section, we discuss intrinsic evaluation or map-level evaluation which evaluates the map directly. While most robotic and embodied AI systems are evaluated based on downstream task success, a few focus on directly evaluating the quality of the maps built during these tasks. However, there is no widely agreed-upon metric suite for evaluating the map representation itself. We next list the various dimensions across which we believe the map can be evaluated and discuss the gaps and future scope for each.
@@ 6.2 Intrinsic evaluation
Accuracy. Map accuracy refers to how accurately the map captures geometric or semantic information when compared against the ground truth. However obtaining the ground truth map can be challenging in most cases. Some works (Georgakis et al., 2022a; Ramakrishnan et al., 2020) evaluate the 2D occupancy maps by using map accuracy to measure how much of that area matches with the ground-truth, and IoU to measure intersection over union between the built and the ground-truth map. While others (Cartillier et al., 2021; Georgakis et al., 2022b) evaluate the accuracy of the built 2D semantic map using semantic segmentation metrics such as pixel-wise labeling accuracy, pixel-based F1 score, IoU score and a contour-based average boundary F1 score. However, there is a lack of standardized metrics that can be used across various map types. For example, IoU is not directly applicable to topological maps. Moreover, while evaluating accuracy for semantic maps that store a fixed set of object categories is straightforward, it is tricky to evaluate semantic maps that store open-vocabulary features. As an alternative, OpenScene (Peng et al., 2023) performs semantic segmentation on their open-vocabulary map for a fixed set of categories and thereafter evaluate it using accuracy and IoU. ConceptGraphs (Gu et al., 2023b), on the other hand, evaluate their open-vocabulary topological map by employing human evaluators on Amazon Mechanical Turk (AMT) to report the scene graph accuracy. OpenLex3D (Kassab et al., 2025) has been recently introduced to evaluate open-vocabulary 3D scene representations, by providing object labels at various granularity or precision. Similar evaluation techniques could benefit intrinsic evaluation in embodied setting as well. The challenge to evaluate modern-day open-vocabulary maps lies in collecting precise ground-truth maps for large real-world environments, which is labor-intensive or infeasible. This limits the use of direct comparisons in order to calculate map accuracy. Moreover, accuracy in dynamic or cluttered environments is also rarely assessed. Both these provide possible future area of research.
@@
Completeness. Map completeness measures how completely a generated map represents an environment, encompassing both geometric and semantic coverage. Geometric coverage refers to the fraction of environment that has been mapped whereas semantic coverage refers to the completeness of the semantic information captured in the map. Completeness is particularly critical for tasks like search and rescue, where a thorough understanding of the surroundings is essential. Moreover, complete geometric and semantic mapping support better decision-making by reducing dependence on incomplete or inaccurate data. The extent of map completeness depends on how thoroughly the robot explores the environment during the downstream task and is closely tied to the ‘stopping criteria’ – a method that determines when to end the exploration process. In embodied AI, exploration typically ends when the task is deemed complete or when a predefined time budget is reached. However, reaching this time budget does not always guarantee that the map is fully complete, making the development of reliable stopping criteria an ongoing challenge in robotics as well as embodied AI research (Placed & Castellanos, 2022; Luperto et al., 2024). Geometric coverage metric has been reported in prior works that performs the task of exploration (Chaplot et al., 2019) in terms of the fraction of the environment explored by the robot. However, a lot of works that tackle other embodied AI tasks fail to report this metric even if exploration is a crucial part of their task (Gervet et al., 2023; Raychaudhuri et al., 2023; Yokoyama et al., 2023). Semantic coverage on the other hand relies on the presence of a detailed ground-truth map with semantic information and could be hard to obtain, as discussed in the previous paragraph ‘Accuracy’. Due to the lack of detailed ground-truth semantic maps, semantic completeness is rarely quantified.
@@
Consistency. Geometric consistency of a map refers to how accurately the spatial structure (distances, angles, and relative positions of structures/objects) of the map represents the physical layout of the environment. Accurate geometry ensures safe and efficient path planning and obstacle avoidance. This is crucial in classical robotics for identifying loop closure, a popular technique to recognize previously visited locations in SLAM, which refines the map and reduces drift. Popular loop closure metrics (Li et al., 2021b; Qian et al., 2022) include Absolute Trajectory Error (ATE) that evaluates the cumulative error, in addition to precision and recall. However, in embodied AI systems, there is no drift due to absence of sensory and actuator noise and hence the generated maps are mostly consistent with the environment. Hence prior embodied AI works do not report geometric consistency metrics for the built map. That said, ensuring geometric consistency will be crucial in future where dynamic moving objects may disrupt the map’s structural fidelity. It can be measured by reporting the Root Mean Square Error (RMSE) from the ground-truth geometry. On the other hand, the semantic consistency of a map refers to the alignment between the semantic information of structures/objects and their physical locations in the environment. Semantic consistency is particularly crucial to remain consistent over time despite changes in perspective, lighting, or environment dynamics when the robot moves around the environment. This could be measured using a temporal accuracy metric that measures how accuracy of the semantic information changes over time. However, none of the prior works measure semantic consistency and may be considered in future research.
Robustness. Evaluating robustness in semantic maps is mostly crucial for assessing their reliability in unpredictable or dynamic environments. A robust map exhibits low uncertainty and high confidence in its semantic information, allowing the robot to adapt to errors, sensor noise, and environmental changes. In SLAM, some study system robustness to measure Absolute Trajectory Error (ATE) and Relative Pose Error (RPE) under noise perturbations (Prokhorov et al., 2019; Yang et al., 2025). On the other hand, since recent approaches use pretrained models, measuring the confidence of the model’s predictions could be beneficial to assess map robustness, with higher confidence indicating a more robust map. Alternatively, model uncertainty can be measured by assessing the variance in model predictions, reflecting epistemic uncertainty in the model, where low uncertainty will indicate a robust map. While some studies in embodied AI (Georgakis et al., 2022a; Raychaudhuri et al., 2025) incorporate uncertainty into task planning, it has not been used as a formal metric in previous research, presenting an area for future exploration. We think that uncertainty-based reasoning in the SLAM and active robotic exploration can provide valuable insights into designing uncertainty-based map robustness metrics. More specifically, joint entropy, expected map information, and uncertainty-aware mapping all relate to reasoning under uncertainty in the robotics literature. Joint entropy quantifies the total uncertainty over both the robot’s trajectory and the map, while expected map information (often expressed as mutual information) captures how much that uncertainty in the map is expected to decrease after a future observation. These are widely used as internal utilities for planning (Stachniss et al., 2005; Blanco et al., 2008a; Georgakis et al., 2022a; Carlone et al., 2013), guiding action selection in active SLAM and exploration by prioritizing trajectories that maximize expected information gain. Uncertainty-aware mapping refers to the practice of explicitly representing and propagating uncertainty within the map itself, often using probabilistic models like occupancy grids with confidence estimates, Gaussian processes, or learned uncertainty models. This enables more robust perception and safer navigation, especially in dynamic or partially observed environments. While joint entropy and expected map information are used to decide where to go next, uncertainty-aware mapping focuses on how we represent what we’ve already seen. However, these are rarely reported as evaluation metrics, as they are task-dependent, abstract, and computationally intensive to calculate precisely.
Table 7: Map evaluation. While extrinsic or task-level evaluation has been extensively studied over the years, there has been little progress in intrinsic evaluation to assess the quality of the built map across various dimensions such as accuracy, completeness, consistency and robustness. In this study, we emphasize the need for a standardized suite of intrinsic metrics to be applied across different map structures and encodings.
Evaluation Dimensions Metrics Task
Extrinsic Utility Success rate, SPL, NE, OSR Navigation (2020; 2020b; 2019b; 2021c; 2024; 2024)
nDTW Instruction following (2020; 2025)
QA accuracy, Nav accuracy Embodied QA (2019a; 2024)
Precision, Recall, F1, ROC Language grounding (2011a; 2012; 2011b)
RPE (translation, rotation) Localization (2019)
Intrinsic Accuracy Map accuracy, IoU, Pixel F1, Boundary F1 Semantic segmentation (2021; 2021)
Completeness Geometric coverage Exploration (2019)
Consistency Precision, Recall, True Positives, ATE Loop closure (2021b; 2022)
Robustness ATE and RPE under noise perturbations Visual SLAM (2019)
6.3 Summary
In this section we summarize the key challenges and the future scope of evaluation (Tab. 7). Extrinsic evaluation metrics have been widely explored in the literature and continue to evolve with the emergence of increasingly complex tasks. However, evaluating the quality of maps (intrinsic) in semantic SLAM and embodied AI continues to be understudied and presents two key challenges. First, collecting accurate, fine-grained, open-vocabulary ground-truth annotations for large, real-world environments is inherently difficult. This limitation makes it particularly challenging to assess important properties of the built maps such as semantic accuracy, consistency, and completeness of the maps. Without reliable ground truth, these evaluations often remain coarse or dataset-specific. Second, there is currently no standardized suite of evaluation metrics that can be consistently applied across different types of map structures (topological, spatial grid, dense geometric, etc.) or across different encoding schemes (explicit and implicit features). Together, these challenges highlight a critical gap in the field and point toward the need for developing scalable, generalizable map evaluation frameworks, offering an important direction for future research.
7 Challenges
Despite significant progress in semantic map building, numerous challenges remain that warrant further attention and improvement. This section outlines key open problems in the field across multiple dimensions.
Efficiency. As semantic maps grow richer by incorporating fine-grained semantic information, they become increasingly data-heavy, posing significant challenges for memory efficiency and compact storage, particularly on robots with constrained hardware. The underlying map structure plays a crucial role in determining storage and computational efficiency. Different representations offer trade-offs among semantic richness, spatial fidelity, and scalability. For example, spatial grids uniformly capture space but are memory-intensive and difficult to scale in large, high-resolution environments. Dense geometric maps provide high spatial fidelity and support per-point semantic features (e.g., color, normals), yet are often redundant and storage-heavy. In contrast, topological maps are lightweight and scalable but lack the geometric detail needed for tasks like manipulation or precise localization. Hybrid topometric maps attempt to balance these aspects by combining the spatial accuracy of grids with the efficiency of topological graphs, though managing and updating both layers increases complexity and computational load. Map encoding methods also impact efficiency. Learned high-dimensional representations (e.g., neural features) offer compressed yet expressive encodings, but they are often harder to interpret or update, and computationally expensive to query. Overall, achieving efficient storage and memory usage in semantic mapping remains an open and active research challenge.
Scalability. As semantic mapping systems are deployed in increasingly large, dynamic, and diverse environments, scalability emerges as a critical challenge. It refers to a system’s ability to accommodate growth in spatial extent (e.g., large physical spaces), semantic complexity (e.g., diverse object categories and scene types), and temporal evolution (e.g., long-term and lifelong operation), while maintaining efficiency, accuracy, and robustness. High-resolution, globally consistent maps require significant computational and memory resources, which grow rapidly with the size and richness of the environment. For example, a home robot navigating multiple cluttered rooms over extended periods must retain detailed spatial layouts and object-level information without exhausting onboard storage or processing capacity. The challenge intensifies in dynamic settings, such as when furniture is rearranged or objects are moved, requiring the map to adapt continuously. Moreover, as maps expand, core operations like loop closure, re-localization and map querying become increasingly expensive in both simulation and real robots. This underscores the need for lightweight, adaptable semantic mapping methods that can scale with environmental complexity while staying efficient on resource-constrained hardware.
Real-time processing. Real-time processing is a critical challenge in semantic map building, especially in applications such as autonomous driving and real-time human-robot interaction, where decisions must be made within strict latency bounds. Semantic mapping in these domains involves constructing maps with semantic and spatial integrity at frame rates fast enough to support safe and responsive behavior. However, this is extremely challenging. For instance, an autonomous vehicle navigating an urban environment must simultaneously detect traffic signals, update the map with moving vehicles or pedestrians, and plan a safe trajectory in milliseconds. Any lag in perception or map update could lead to disastrous decisions. Balancing the trade-off between semantic richness and processing speed, often on memory-constrained hardware, remains an open research challenge. Advances in model compression, edge computing, and efficient data representations are being explored, but ensuring high-fidelity semantic understanding under tight latency requirements continues to be a major hurdle for real-world deployment.
Noise and uncertainty. Handling noise and uncertainty is a fundamental challenge in semantic map building, particularly when operating in real-world environments characterized by sensor imperfections, dynamic changes and ambiguous semantics. Noise from cameras, LiDARs, or depth sensors can lead to inaccurate observations, while errors in object detection, segmentation, or classification models may introduce semantic inconsistencies in the map. For example, a robot navigating a cluttered indoor environment might repeatedly misclassify similar-looking objects (mistaking a stool for a chair) or fail to detect partially occluded objects or objects in poor lighting. Over time, these errors can compound, leading to semantic drift or contradictions in the map. Furthermore, the uncertainty in sensor data and learned representations is often not explicitly modeled or propagated, which limits the system’s ability to reason about the confidence of its predictions. This becomes especially critical in safety applications where overconfident yet incorrect map entries can mislead planning and decision-making. While some approaches use Bayesian filters or probabilistic representations to quantify uncertainty in robotics (Blanco et al., 2008a; Georgakis et al., 2022b), these are computationally expensive and may not scale well. Moreover, semantic mapping methods in embodied AI often assume noiseless sensors in simulated environments, and thus do not transfer well to real robots. Thus, building robust semantic maps that explicitly model and manage uncertainty remains a key open problem in both robotics and embodied AI.
Multi-modal fusion. Recent semantic map building approaches increasingly rely on fusing diverse sensory inputs, including vision, depth, audio, natural language, speech etc., to construct rich interpretable representations of the environment. However, robust multi-modal fusion remains a core challenge, with alignment and integration between the modalities a non-trivial problem. For instance, visual inputs provide spatial and appearance cues, while natural language and speech often provide high-level semantic cues that may be ambiguous, indirect or context-dependent (e.g., ‘go to the room where the baby is sleeping’). Moreover such grounding is often compounded by noisy or incomplete observations. Audio cues such as footsteps, conversations or object sounds offer valuable environmental context but are transient and hard to spatially localize or persistently encode. Moreover, multi-modal fusion needs to be done in a way that supports downstream reasoning and generalization in unseen environments. For instance, enabling agents to query the semantic map for complex spatial reasoning (e.g. ‘the cup behind the fruit bowl’, ‘the vase between the kettle and the dish’) and affordances (e.g. ‘where can I sit’, ‘which objects are fragile’). The emergence of multi-modal foundation models (Alayrac et al., 2022; Driess et al., 2023; OpenAI, 2023) offers new promise in addressing some of these challenges. These models are capable of aligning and reasoning across modalities, enabling semantic parsing of complex instructions, visual understanding and basic spatial reasoning. However, they are not trained explicitly for map construction or long-term memory integration, limiting their direct applicability in real-time embodied settings. Moreover, deploying such large models on robot hardware poses computational and energy constraints, particularly in latency-sensitive domains like autonomous driving. Thus the challenge lies in building efficient multi-modal fusion strategies that also support reliable and flexible querying, making this an open research problem.
Lifelong learning. Lifelong learning poses significant challenges for semantic mapping, particularly in real-world applications where robots are expected to operate continuously over extended periods in dynamic environments. Unlike static maps that assume a fixed world, real-world settings are inherently non-stationary (furniture is rearranged, objects are moved, and scenes evolve with time and seasons). To support lifelong learning and long-term autonomy, semantic maps must be capable of continuous adaptation, i.e. updating outdated information, integrating new observations and distinguishing between transient and persistent changes. This requires temporal reasoning, mechanisms to prevent catastrophic forgetting and the ability to resolve conflicting data collected across different times or viewpoints. The importance of long-term autonomy lies in its potential for practical deployment, i.e. robots that can adapt to changing homes, warehouses or outdoor environments without manual reconfiguration. It is a foundational requirement for truly intelligent agents that not only act within their world but grow and evolve alongside it, building richer and more useful semantic representations over time. However, most current systems are not yet equipped with robust mechanisms for continuous semantic learning and remain limited in handling evolving, open-world environments.
Standardized evaluation framework. A significant challenge in semantic map building lies in the absence of standardized evaluation frameworks that can consistently benchmark performance across diverse tasks, environments and map representations, as we discuss at length in Sec. 6.2. While the community has focused more on task-specific extrinsic evaluation, limited progress have been made towards developing intrinsic measures to evaluate the map quality. This lack of standardized intrinsic evaluation hinders progress, as it becomes unclear whether improvements in downstream task success stem from better semantic understanding or other factors like better control policies. Developing general, interpretable and task-agnostic evaluation metrics is crucial, not only to track progress but also to facilitate cross-domain generalization. As open-vocabulary multi-modal queryable general-purpose semantic maps become increasingly central to reasoning in embodied AI and real-world robots, creating a unified benchmark suite that reflects both structural and semantic fidelity is essential for advancing the field.
Summary. Several critical challenges remain unaddressed in semantic mapping approaches across various dimensions. Key open problems include building efficient, scalable and real-time maps, which are robust to noise and uncertainty. Building multi-modal maps, that are reliable and flexible, too remains an open challenge. Maps that are able to handle lifelong learning also remain under-explored. Moreover, lack of a standardized evaluation framework require further research and innovation.